turn 13.1% into a fraction
Answer by jim_thompson5910(35256)

13.1% = 13.1/100 = 131/1000


So 13.1% as a fraction is 131/1000

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim